# Rhetorical XAI: Explaining AI’s Benefits as well as its Use via Rhetorical Design

### v1

@@ Abstract.
This paper explores potential benefits of incorporating Rhetorical Design into the design of Explainable Artificial Intelligence (XAI) systems. While XAI is traditionally framed around explaining individual predictions or overall system behavior, explanations also function as a form of argumentation, shaping how users evaluate system perceived usefulness, credibility, and foster appropriate trust. Rhetorical Design offers a useful framework to analyze the communicative role of explanations between AI systems and users, focusing on: (1) logical reasoning conveyed through different types of explanations, (2) credibility projected by the system and its developers, and (3) emotional resonance elicited in users. Together, these rhetorical appeals help us understand how explanations influence user perceptions and facilitate AI adoption. This paper synthesizes design strategies from prior XAI work that align with these three rhetorical appeals and highlights both opportunities and challenges of integrating rhetorical design into XAI design.
@@ 1. Introduction
Different stakeholders leverage explanations to pursue distinct goals (Kim et al., 2024; Hoffman et al., 2021; Preece et al., 2018; Tomsett et al., 2018; Mohseni et al., 2021). For example, developers employ explanations to improve predictive performance by debugging model errors and biases (Hohman et al., 2018; Molnar, 2020). Ethicists and policymakers rely on explanations to assess the safety and accountability of AI systems (Busuioc, 2021). HCI researchers investigate how explanations can foster user appropriate trust (Zhang et al., 2020; Lai et al., 2023; Vasconcelos et al., 2023) and support complementary human-AI teaming (Bansal et al., 2020). Stakeholders with diverse objectives tend to evaluate explanations through different lenses, shaped by their specific roles and needs. In particular, end-users tend to judge explanations alongside more pragmatic and affective dimensions such as utility, trustworthiness, and alignment with personal values (Vera Liao et al., 2022). Therefore, a faithful and logically accurate explanation that neglects human-centered considerations may be disregarded altogether.
In this paper, we argue that designing explanations for effective adoption motivates expanding the scope of XAI beyond explaining how AI works towards articulating why AI merits use. An AI system is just one of many possible solutions to a user problem and AI explanations can serve a rhetorical function by communicating why such a system is beneficial, credible, and appropriate in a given context. Applications might include supporting critical evaluations of AI decisions (Wang and Yin, 2022; Bansal et al., 2020; Nguyen et al., 2018), assessing system credibility (Binns et al., 2018; Dodge et al., 2019), and calibrating appropriate user trust (Markus et al., 2021; Liao and Sundar, 2022).
@@
Our paper offers both conceptual and practical contributions to the design of XAI systems. Conceptually, we propose a rhetorical framework for XAI that expands the traditional focus on explaining how AI works to include why AI merits use. Practically, guided by this framework, we analyze existing design strategies through the three classical rhetorical appeals (logos, ethos, and pathos), highlighting how explanations can be designed to be more logically convincing, credibly appealing, and emotionally engaging. By critically assessing the strengths and limitations of these strategies, we also discuss their broader implications for future XAI research and design.
@@ 2.1. XAI Technical Concepts
Building on this perspective, the following sections briefly introduce primary XAI design paradigms including user-centered (Section 2.2), stakeholder-centered (Section 2.3), and sociotechnical approaches to explanation (Section 2.4).
@@ 2.5. Research Focus of this Study
Thus far, Section 2 has outlined key technical concepts in XAI, covering commonly used terminologies and taxonomies (2.1), as well as design paradigms of human-centered XAI, including user-centered (2.2), stakeholder-centered (2.3), and sociotechnical perspectives (2.4).
In this work, we expand the design paradigms of human-centered XAI to incorporate the rhetorical dimensions of user experience. While traditional XAI prioritizes functional utility and cognitive human-AI alignment across different stakeholder objectives, we argue that pragmatic and affective responses equally shape understanding and adoption of AI (Dewey, 2005; Buchanan, 1985a, 2001a). For example, when interacting with XAI systems, users frequently grapple with questions such as: Will this explanation help me achieve my goal better? Can I trust its rationale? Does it align with my ethical principles? These questions highlight that technical fidelity and cognitive alignment alone are insufficient; explanations must also resonate with users’ expectations, emotions, and values to be meaningful and persuasive.
Our reframing positions AI explanations not only as informative responses but also as rhetorical artifacts that mediate the relationship between AI systems and end-users. In other words, AI is just one of many possible solutions to a user problem. AI explanations thus must serve a persuasive and contextualizing role by communicating why an AI system is beneficial, credible, and appropriate in a given context. From this perspective, a rhetorical XAI design can more effectively support users in evaluating automated decisions (Wang and Yin, 2022; Bansal et al., 2020; Nguyen et al., 2018), assessing system credibility (Binns et al., 2018; Dodge et al., 2019), and calibrating appropriate user trust (Markus et al., 2021; Liao and Sundar, 2022).
In this paper, we investigate the argumentative role of AI explanations in mediating user adoption of AI systems and we synthesize how prior research has employed various design strategies to fulfill this rhetorical function. We draw on the concept of Rhetorical Design, a longstanding product and graphic design practice that examines how form, structure, and presentation influence user adoption. In the next section, we introduce the theoretical foundations and design practices of rhetorical design in HCI and describe how we adapt this lens to examine and inform explanation design.
@@ 3.3. A Rhetorical Framework to Characterize Explanation Design
Building on this conceptual adaptation, we argue that rhetorical design can serve two distinct objectives in XAI: (1) supporting traditional AI explanations that clarify How AI Works, and (2) designing explanations that communicate Why AI Merits Use to encourage adoption. We examine each objective through the three rhetorical dimensions: logical reasoning, credibility, and emotional resonance. In Table 1, we outline each rhetorical appeal as a coding inquiry across the two goals, which guided our open and axial coding process detailed in the next section.
@@ 4. Method
We conducted a narrative review guided by coding inquiries specified in Table 1. Unlike a systematic review, which follows established protocols such as PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) (Moher et al., 1996), we chose a narrative review to allow for greater flexibility in selecting, interpreting, and synthesizing diverse design strategies across different XAI contexts (Baumeister and Leary, 1997; Ferrari, 2015). Nevertheless, potential author bias remains an inherent limitation of this approach (see limitations in Section 6.3). To mitigate this, we employed a structured three-step process involving both open and axial coding.
@@
Additionally, we contextualized quantitative measures with reported qualitative feedback. We provide our reviewed articles and corresponding coding schemes in Appendix B. For each paper, we specified the primary XAI goal, identified design strategies, and mapped corresponding quantitative and qualitative feedback to appropriate rhetorical appeals.
@@ 5.1.1. Isolated Presentation
Different types of AI explanations inherently embody distinct forms of human logical reasoning. Wang et al. (2019) conceptually mapped how various machine learning models reason and generate explanations to corresponding forms of human decision-making. For example, they argued that Bayesian probability represents inductive reasoning, a bottom-up approach that uses prior intuitions to draw broader conclusions. Case-based models encapsulate analogical reasoning, which solves new problems by synthesizing solutions from previously solved similar issues. Rule-based models illustrate deductive reasoning, a top-down approach that begins with general principles and operationalizes them to specific cases.
@@
A dominant design strategy employed in prior XAI studies is isolated presentation, where different types of explanations are presented individually for A/B testing. Specifically, these studies examined the extent to which distinct logical structures embodied in different explanations (e.g., features, examples, rules, or concepts) helped to improve human-AI complementary performance (Bansal et al., 2020; Gajos, 2024; Nguyen et al., 2018) or promote appropriate reliance on AI predictions (Schemmer et al., 2023; Yang et al., 2020; Zhang et al., 2020; Liu et al., 2021; Vasconcelos et al., 2023). For example, in comparing feature-based and example-based explanations, Chen et al. (2023b) observed that the latter elicited more critical reflections on potential misalignments between human intuition and machine reasoning. Moreover, van der Waa et al. (2021) demonstrated that rule-based explanation showed a small positive effect on (global) system understanding compared to example-based explanation. Nevertheless, both types of explanations could lead to confirmation bias (Bertrand et al., 2022), where users tended to over-rely on AI predictions erroneously.
@@ 5.1.2. Adaptive and Mechanistic Presentation
For system-triggered deliveries, explanations are automatically adjusted in response to internal (invisible to end-users) indicators. For example, some XAI systems present explanations only when AI confidence is low (Bansal et al., 2020) or dynamically adjust explanations based on real-time accuracy updates or error boundaries (Bansal et al., 2019b, a; Wang and Yin, 2023). Others adapt explanations based on human learning progress (Wambsganss et al., 2020) or show explanations only when the AI’s predicted correctness is estimated to exceed that of the user (Ma et al., 2023).
@@
Overall, both adaptive and mechanistic presentations offer unique benefits to improve user understanding and support human-AI teaming. Adaptive presentations enhance relevance by tailoring explanations to user goals, expertise, or decision context, reducing cognitive overload and aligning responses with moment-to-moment needs. Mechanistic presentations offer breadth and transparency by juxtaposing multiple explanation types in a unified interface, allowing users to cross-reference local and global reasoning or compare factual and counterfactual views. Compared to isolated presentations, interactivity is a crucial differentiating factor, allowing logical reasoning processes to unfold dynamically in response to user actions or system feedback.
@@ 5.1.3. Graphical Synergy
Regardless of whether explanations are presented in isolated, adaptive, or mechanistic formats, different explanation types often coexist unilaterally without being cohesively integrated. As Wang et al. (2019) noted, this fragmentation arises because the underlying explanatory algorithms rely on fundamentally different logical bases. Nevertheless, emerging work has shown promise in organically merging diverse explanation types into a visually coherent representation, an approach we term graphical synergy.
@@ 5.1.4. Explanations as Recommenders
Traditionally, XAI work focused on explaining how systems work by helping users to assess, replicate, or predict AI decisions. Emerging work now shifts toward explaining why AI systems merit use, especially by supporting users in solving real-world problems beyond simple classification. In these contexts, explanations act as recommenders, helping users understand what actions to take and why those actions are beneficial.
For example, Wambsganss et al. (2020) developed an adaptive AI-assisted learning system to assist students in writing persuasive arguments. To facilitate revisions, the system visualized argumentative structure, evaluated writing quality, and provided natural language feedback alongside highlighted areas for improvement. Students expressed a significantly greater preference for using the new system than traditional reference-based writing tools. Nilforoshan and Wu (2018) implemented a similar application using feature-based explanations to improve reviewer responses on social media.
@@
Findings from Wang et al. (2023b) highlight a key point that aligns with our framework (outlined in Section 3.3): effective use of an AI system does not necessarily require a deep understanding of system behaviors. For example, in their user study, participants rated “ease of understanding” (4.90) and “help with understanding the model” (5.17) lower than their “intention to reuse the system for future loan applications” (5.34) and “the system’s usefulness in identifying ways to improve their loan applications” (5.61), based on a 7-point Likert scale. Although observed differences were not statistically validated, this finding suggests that a deep understanding of model behaviors is not a prerequisite for effectively using the AI for algorithmic recourse. This reinforces our position on XAI’s rhetorical role: explanations should reveal how AI works and why an AI merits use for practical utility.

### v2

@@ Abstract.
We explore potential benefits of incorporating Rhetorical Design into the design of Explainable Artificial Intelligence (XAI) systems. While XAI is traditionally framed around explaining individual predictions or overall system behavior, explanations also function as a form of argumentation, shaping how users evaluate system perceived usefulness, credibility, and foster appropriate trust. Rhetorical Design offers a useful framework to analyze the communicative role of explanations between AI systems, and end-users, focusing on: (1) logical reasoning conveyed through different types of explanations, (2) credibility projected by the system and its developers, and (3) emotional resonance elicited in users. Together, these rhetorical appeals help us understand how explanations influence people perceptions and facilitate AI adoption across and within different collaborative and social contexts. This paper synthesizes design strategies from prior XAI work that align with these three rhetorical appeals and highlights both opportunities and challenges of integrating rhetorical design into XAI design.
Explainable AI, Rhetorical Design
@@ 1. Introduction
Different stakeholders leverage explanations to pursue distinct goals (Kim et al., 2024; Hoffman et al., 2021; Preece et al., 2018; Tomsett et al., 2018; Mohseni et al., 2021). For example, developers employ explanations to improve predictive performance by debugging model errors and biases (Hohman et al., 2018; Molnar, 2020). Ethicists and policymakers rely on explanations to assess the safety and accountability of AI systems (Busuioc, 2021). HCI researchers investigate how explanations can foster user appropriate trust (Zhang et al., 2020; Lai et al., 2023; Vasconcelos et al., 2023) and support complementary human-AI teaming (Bansal et al., 2021). Stakeholders with diverse objectives tend to evaluate explanations through different lenses, shaped by their specific roles and needs. In particular, end-users tend to judge explanations alongside more pragmatic and affective dimensions such as utility, trustworthiness, and alignment with personal values (Vera Liao et al., 2022). Therefore, a faithful and logically accurate explanation that neglects human-centered considerations may be completely disregarded.
In this paper, we argue that designing explanations for effective adoption motivates expanding the scope of XAI beyond explaining how AI works towards also articulating why AI merits use. An AI system is just one of many possible solutions to a user problem and AI explanations can serve a rhetorical function by communicating why such a system is beneficial, credible, and appropriate in a given context. Applications might include supporting critical evaluations of AI decisions (Wang and Yin, 2022; Bansal et al., 2021; Nguyen et al., 2018), assessing system credibility (Binns et al., 2018; Dodge et al., 2019), and calibrating appropriate user trust (Markus et al., 2021; Liao and Sundar, 2022).
@@
Our paper offers both conceptual and practical contributions to the design of XAI systems. Conceptually, we propose a rhetorical framework for XAI that expands the traditional focus on explaining how AI works to include why AI merits use. Practically, guided by this framework, we analyze existing design strategies through the three classical rhetorical appeals (logos, ethos, and pathos), highlighting how explanations can be designed to be more logically convincing, credibly appealing, and emotionally engaging. By integrating these communicative functions, we show how AI explanations can extend beyond technical transparency to become richer rhetorical artifacts that help diverse stakeholders interpret AI outputs, negotiate responsibilities, build trust, and coordinate decisions in collaborative and social contexts. By critically assessing the strengths and limitations of these strategies, we also discuss their broader implications for future XAI research and design.
@@ 2.1. XAI Technical Concepts
Building on this perspective, the following sections briefly introduce XAI design paradigms including user-centered (Section 2.2), stakeholder-centered (Section 2.3), and sociotechnical approaches to explanation (Section 2.4).
@@ 2.5. Research Focus of this Study
Thus far, Section 2 has outlined key technical concepts in XAI, covering commonly used terminologies and taxonomies (2.1), as well as design paradigms of human-centered XAI, including user-centered (2.2), stakeholder-centered (2.3), and sociotechnical perspectives (2.4). This section aims to differentiate our work by conceptualizing AI explanations as rhetorical artifacts that mediate communication between designers and end-users, beyond merely justifying outputs or revealing flaws.
While user-centered XAI (Section 2.2) prioritizes functional utility and cognitive human-AI alignment across different user objectives, we argue that pragmatic and affective responses equally shape understanding and adoption of AI (Dewey, 2005; Buchanan, 1985a, 2001a). For example, when interacting with XAI systems, users frequently grapple with questions such as: Will this explanation help me achieve my goal better? Can I trust its rationale? Does it align with my ethical principles of using AI? These questions highlight that technical fidelity and cognitive alignment alone are insufficient; explanations must also resonate with user expectations, emotions, and values to be meaningful and persuasive.
More importantly, while the stakeholder-centered framework (Section 2.3) broadens the scope of XAI to include the goals and values of broader stakeholder groups, and socio-technical focus (Section 2.4) highlights the practical challenges of implementing XAI in an organizational setting, both still frame AI explanations primarily as tools either to justify AI results or to expose their flaws. We argue that this framing neglects the role of explanation as a communicative act in AI use, which is shaped by design and situated within a social context, inviting diverse interpretations around why AI appears logically sound, credibly useful, and emotionally appealing. When multiple stakeholders engage with explanations in collaborative settings, it is important to understand how different explanation designs shape their interpretations. These interpretations, in turn, influence perceptions of AI transparency, the negotiation of responsibilities, the development of trust, and the coordination of final AI adoption in practice.
Our reframing positions AI explanations not only as informative responses but also as rhetorical artifacts that mediate the relationship between AI systems and end-users. In other words, AI is just one of many possible solutions to a user problem. AI explanations thus must serve a persuasive and contextualizing role by communicating why an AI system is beneficial, credible, and appropriate in different contexts. From this perspective, a rhetorical XAI design can more effectively support stakeholders in evaluating automated decisions (Wang and Yin, 2022; Bansal et al., 2021; Nguyen et al., 2018), assessing system credibility (Binns et al., 2018; Dodge et al., 2019), and calibrating appropriate user trust (Markus et al., 2021; Liao and Sundar, 2022).
In this paper, we investigate the argumentative role of AI explanations in mediating user adoption of AI systems, and we synthesize how prior research has employed various design strategies to fulfill this rhetorical function. We draw on the concept of Rhetorical Design, a longstanding product and graphic design practice that examines how form, structure, and presentation influence user adoption. In the next section, we introduce the theoretical foundations of rhetorical design in HCI and describe how we adapt this lens to examine and inform explanation design.
@@ 3.3. A Rhetorical Framework to Characterize Explanation Design
Building on this conceptual adaptation, we argue that rhetorical design can serve two complementary objectives in XAI: (1) supporting traditional AI explanations that clarify How AI Works, and (2) designing explanations that communicate Why AI Merits Use to guide or inform adoption. Our goal in foregrounding this distinction is not to enforce a rigid dichotomy, but to highlight the nuanced differences between these design objectives and propose corresponding strategies. More importantly, we hope that clarifying both objectives and keeping them in mind during the design process can enhance the value of explanations, suggesting which types of explanations to use and the contexts where they are appropriate or inappropriate regarding the technology affordance, task requirements, and the stakes involved (we discuss this further in Section 6.1).
We examine each objective through the three rhetorical dimensions: logical reasoning, credibility, and emotional resonance. In Table 1, we outline each rhetorical appeal as a coding inquiry across the two goals, which guided our open and axial coding process detailed in the next section. While Table 1 contrasts the two explanatory goals to clarify our coding process, this framing is intended to support, not discourage, explanation strategies that span both objectives.
@@ 4. Method
We conducted a narrative review guided by coding inquiries specified in Table 1. Unlike a systematic review, which follows established protocols such as PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) (Moher et al., 1996), we chose a narrative review to allow for greater flexibility in selecting, interpreting, and synthesizing diverse design strategies across different XAI contexts (Baumeister and Leary, 1997; Ferrari, 2015). We detail the review process in Appendix B, including search terms, databases, inclusion/exclusion criteria, and paper grouping strategies. Nevertheless, potential author bias remains an inherent limitation of this approach (see limitations in Section 6.3). To mitigate this, we employed a structured three-step process involving both open and axial coding.
@@
Additionally, we contextualized quantitative measures with reported qualitative feedback. We present the reviewed articles and their corresponding coding schemes in Table LABEL:tab:coding_scheme. For each paper, we specified the primary XAI goal, identified design strategies, and mapped both quantitative and qualitative feedback to the appropriate rhetorical appeals. Each code assigned to an excerpt from the reviewed papers indicates the design intent it represents and the rhetorical appeal it conveys.
@@ 5.1.1. Isolated Presentation
Different types of AI explanations (see Appendix A) inherently embody distinct forms of human logical reasoning. Wang et al. (2019) conceptually mapped how various machine learning models reason and generate explanations to corresponding forms of human decision-making. For example, they argued that Bayesian probability represents inductive reasoning, a bottom-up approach that uses prior intuitions to draw broader conclusions. Case-based models encapsulate analogical reasoning, which solves new problems by synthesizing solutions from previously solved similar issues. Rule-based models illustrate deductive reasoning, a top-down approach that begins with general principles and operationalizes them to specific cases.
@@
A dominant design strategy employed in prior XAI studies is isolated presentation, where different types of explanations are presented individually for A/B testing. Specifically, these studies examined the extent to which distinct logical structures embodied in different explanations (e.g., features, examples, rules, or concepts) helped to improve human-AI complementary performance (Bansal et al., 2021; Gajos, 2024; Nguyen et al., 2018) or promote appropriate reliance on AI predictions (Schemmer et al., 2023; Yang et al., 2020; Zhang et al., 2020; Liu et al., 2021; Vasconcelos et al., 2023). For example, in comparing feature-based and example-based explanations, Chen et al. (2023b) observed that the latter elicited more critical reflections on potential misalignments between human intuition and machine reasoning. Moreover, van der Waa et al. (2021) demonstrated that rule-based explanation showed a small positive effect on (global) system understanding compared to example-based explanation. Nevertheless, both types of explanations could lead to confirmation bias (Bertrand et al., 2022), where users tended to over-rely on AI predictions erroneously.
@@
Based on extensive prior research examining the user effect of various explanations, presenting isolated explanations may be most suitable in early-stage design and evaluation settings, where the goal is to investigate how different explanation formats align with the needs (Liao et al., 2020) and preference of a target stakeholder group (Dhanorkar et al., 2021; Panigutti et al., 2023; Tonekaboni et al., 2019). This approach is also effective for formulating user mental models in controlled experiments, such as identifying the error boundaries of human-AI decision-making (Bansal et al., 2019a, b). Notably, isolated explanations provide a clear basis for comparison before progressing to more advance designs, such as adaptive and mechanistic presentation, or graphical synergy, which reveal AI behaviors in greater complexity (as described in the following sections).
@@ 5.1.2. Adaptive and Mechanistic Presentation
For system-triggered deliveries, explanations are automatically adjusted in response to internal (invisible to end-users) indicators. For example, some XAI systems present explanations only when AI confidence is low (Bansal et al., 2021) or dynamically adjust explanations based on real-time accuracy updates or error boundaries (Bansal et al., 2019b, a; Wang and Yin, 2023). Others adapt explanations based on human learning progress (Wambsganss et al., 2020) or show explanations only when the AI’s predicted correctness is estimated to exceed that of the user (Ma et al., 2023).
@@
Overall, both adaptive and mechanistic presentations offer unique benefits to improve user understanding and support human-AI teaming. Adaptive presentations enhance relevance by tailoring explanations to user goals, expertise, or decision context, reducing cognitive overload and aligning responses with moment-to-moment needs. Thus, adaptive explanations are particularly valuable in dynamic or iterative workflows, where user needs may evolve (Suresh et al., 2021). Mechanistic presentations offer breadth and transparency by juxtaposing multiple explanation types in a unified interface, allowing users to cross-reference local and global reasoning or compare factual and counterfactual views. As a result, they are more effective in diagnostic or investigative tasks, where users benefit from cross-checking different explanatory lenses to reach more robust conclusions (Miller, 2023a). Compared to isolated presentations, interactivity is a crucial differentiating factor for adaptive and mechanistic presentations, allowing logical reasoning processes to unfold dynamically in response to user actions or system feedback.
@@ 5.1.3. Graphical Synergy
Regardless of whether explanations are presented in isolated, adaptive, or mechanistic formats, different explanation types often coexist unilaterally without being cohesively integrated. As Wang et al. (2019) noted, this fragmentation arises because the underlying explanatory algorithms rely on fundamentally different logical bases. Nevertheless, emerging work has shown promise in organically merging diverse explanation types into a visually coherent representation, an approach we term graphical synergy. Due to its visual nature, graphical synergy can be effective for pattern recognition tasks, where synthesizing complementary explanation types helps users grasp subtle relationships that no single explanation could convey.
@@ 5.1.4. Explanations as Recommenders
Traditionally, XAI work focused on explaining how systems work by helping users to assess, replicate, or predict AI decisions. Emerging work now shifts toward explaining why AI systems merit use, particularly by supporting users in solving real-world problems beyond simple classification. In these contexts, explanations act as recommenders, helping users understand what actions to take and why those actions are beneficial.
For example, Wambsganss et al. (2020) developed an adaptive AI-assisted learning system to help students write persuasive arguments. To facilitate revisions, the system visualized argumentative structure, evaluated writing quality, and provided natural language feedback alongside highlighted areas for improvement. Students expressed a significantly greater preference for using the new system than traditional reference-based writing tools. Nilforoshan and Wu (2018) implemented a similar application using feature-based explanations to improve reviewer responses on social media.
@@
Findings from Wang et al. (2023b) highlight a key point that aligns with our framework (outlined in Section 3.3): effective use of an AI system does not necessarily require a deep understanding of system behaviors. For example, in their user study, participants rated “ease of understanding” (4.90) and “help with understanding the model” (5.17) lower than their “intention to reuse the system for future loan applications” (5.34) and “the system’s usefulness in identifying ways to improve their loan applications” (5.61), based on a 7-point Likert scale. Although observed differences were not statistically validated, this finding suggests that a deep understanding of model behaviors is not a prerequisite for effectively using the AI for algorithmic recourse.
To summarize, explanations as recommenders are particularly beneficial in scenarios where users seek actionable guidance. By positioning AI as a collaborative advisor, this strategy reinforces our position on XAI’s rhetorical role: explanations should reveal how AI works and why an AI merits use for practical utility.